OpenStudy (anonymous):

Find the first derivative of each of the following function and leave your answer in its simplest form: i. y=2x^-5 + 3(sqrt x) - 8/7 dy/dx = -10x^-6 + 3/2 x ^(-1/2) ??

OpenStudy (anonymous):

ii. r=((t³-2)/(4t⁵+19)) U=(t^3)-2 v= 4t^5 +19 du/dx=3t^2 dv/dx=20t^4 (v(du/dx)- u(dv/dx))/v^2 (4t^5 +19)(3t^2)-((t^3)-2 )(20t^4) ------------------------------ ( 4t^5 +19)^2 12t^7+57t^2 - 20t^7 + 40t^4 ---------------------------- ( 4t^5 +19)^2 -8t^7+40t^4-57t^2 ------------------ ( 4t^5 +19)^2 t^2(-8t^5+40t^2-57) ------------------- ( 4t^5 +19)^2

OpenStudy (anonymous):

is my answers correct???